OpenStudy (anonymous):

what would the graph of 10/x look like? A straight line a parabola a dotted line none of the above

OpenStudy (anonymous):

@e.cociuba @Ell @erendida @Ebbaybbay @robtobey @radar @redred413 @RolyPoly @RANE @rachelle_known1 @TaeDoingItBig @trasytas @thomaster

OpenStudy (jkristia):

try with a couple of sample points on both sides of x = 0. E.g -5, -2, -1, 1, 2, 5 Is the function defined at x = 0?

OpenStudy (anonymous):

I don think it is

OpenStudy (jkristia):

Correct, so if it is not defined at x=0, then you have a vertical asymptote.

OpenStudy (jkristia):

did you try plot the sample points ?

OpenStudy (anonymous):

I don't have graph paper with me but ima try it in paint

OpenStudy (anonymous):

a parabola I think

OpenStudy (jkristia):

no, not a parabole. You done need graph paper, just make draw x/y coordinate on a blank paper, add 5 units in each direction, fill in the sample points f(-5) = 10/-5 = -2 f(-2) = 10 / -2 = -5 f(-1) = 10 / -1 = -10 f(1) = 10 / 1 = ? f(2) = 10 / 2 = ? f(5) = 10 / 5 = ?

OpenStudy (anonymous):

uh sorry I don't understand?

OpenStudy (jkristia):

 Not the best, but it gives an idea of what the graph looks like, with just a few sample points on each side of x = 0